At 6am Allison left Seattle and headed south at 20 mph. At 8am Michael left from the same place and headed south on the same road. If Michael was 50 miles ahead by 1pm on the same day, what was his speed?
Answer by richwmiller(17219)

2 hours at 20 mph =40 miles
40+50=90
90/5=18
20+18=38 mph Michael's speed
check
20*7=140
38*5=190
190-140=50 miles ahead